### v1

# Learning Invariant Causal Mechanism from Vision-Language Models

Abstract
Pre-trained large-scale models have become a major research focus, but their effectiveness is limited in real-world applications due to diverse data distributions. In contrast, humans excel at decision-making across various domains by learning reusable knowledge that remains invariant despite environmental changes in a complex world. Although CLIP, as a successful vision-language pre-trained model, demonstrates remarkable performance in various visual downstream tasks, our experiments reveal unsatisfactory results in specific domains. Our further analysis with causal inference exposes the current CLIP model’s inability to capture the invariant causal mechanisms across domains, attributed to its deficiency in identifying latent factors generating the data. To address this, we propose the Invariant Causal Mechanism of CLIP (CLIP-ICM), an algorithm designed to provably identify invariant latent factors with the aid of interventional data, and perform accurate prediction on various domains. Theoretical analysis demonstrates that our method has a lower generalization bound in out-of-distribution (OOD) scenarios. Experimental results showcase the outstanding performance of CLIP-ICM.
Machine Learning, ICML
1 Introduction
Pre-trained large-scale models have garnered sustained attention as a prominent research direction (Chen et al., 2020; He et al., 2022; Radford et al., 2021). The primary objective of these models is to provide a unified representation for various downstream tasks (Bengio et al., 2014).
In real-world applications, these models are supposed to be flexibly adapted to more diverse scenarios, encompassing data distribution different from those encountered during training (Torralba & Efros, 2011; Fang et al., 2013; Beery et al., 2018; Gulrajani & Lopez-Paz, 2020; Abbe et al., 2023). Consequently, the generalization ability of pre-train models is restricted, leading to the formation of out-of-distribution (OOD) problems.
Figure 1: The zero-shot prediction of CLIP vs our method.
Existing approaches including (Ganin et al., 2016; Li et al., 2018; Arjovsky et al., 2020; Sagawa* et al., 2019) have been designed to mitigate the challenges associated with OOD scenarios. While these methods have achieved a certain degree of success in addressing related issues, the OOD problem still persists and demands urgent resolution (Gulrajani & Lopez-Paz, 2020). Providentially, vision-language representation learning presents promising alternatives (Frome et al., 2013; Socher et al., 2013). Vary from traditional label-based visual pre-train models and unsupervised pre-train models, vision-language representation learning enriches representations by aligning images with natural language descriptions, offering a wealth of semantic information. Among the most widely adopted vision-language pre-train models, Contrastive Language-Image Pretraining (CLIP) (Radford et al., 2021) stands out for its ability to seamlessly transfer to various visual downstream tasks in a zero-shot manner, yielding commendable performance.
To assess the efficacy of CLIP in addressing OOD issues, we investigate CLIP’s classification performance across different domains within the VLCS (Fang et al., 2013) dataset. However, experimental results reveal inconsistent performance of CLIP when applied to various domains. In particular, CLIP shows poor performance when faced with the OOD problem in zero-shot scenarios. In other words, the representations generated by CLIP do not capture all the invariant features shared across domains. The human cognitive process involves highly abstract representations. In the process of understanding the world, humans tend to encapsulate the mechanisms of the physical world as a collection of causal relationships (Schölkopf et al., 2021). Humans are capable of making robust judgments across various environments based on features that remain invariant across different domains (Peters et al., 2015; Arjovsky et al., 2020; Ahuja et al., 2022a). However, CLIP fails to gain the capacity to make this kind of judgment.
To further validate our results, we construct a structural causal model (SCM) (Pearl, 2009) for the prediction process of CLIP in OOD scenarios. Additionally, we demonstrate that the current prediction process fails to capture invariant causal mechanisms. Furthermore, we identify and substantiate the sufficient and necessary conditions for capturing invariant causal mechanisms. Based on these conditions, we propose CLIP-ICM which could effectively capture the invariant causal mechanisms across various domains. Subsequently, we conduct a generalization analysis of the model based on Rademacher complexity, demonstrating that CLIP-ICM has a lower generalization bound compared to CLIP. This observation implies a superior generalization ability of our method. Experimental results on commonly used OOD datasets corroborate the outstanding effectiveness of our approach in achieving superior performance under zero-shot and linear-probe conditions.
Our contributions could be summarized as follows:
  • We discover that CLIP in OOD scenarios always suffers from the poor performance problem. Then, we establish a CLIP-related SCM from a causal perspective, demonstrating its inability to capture invariant causal mechanisms. Subsequently, we identify the sufficient and necessary conditions for capturing invariant causal mechanisms.
  • We give a theoretical analysis of identifying invariant latent factors in feature representations under a specific condition, and propose a method called CLIP-ICM to satisfy that specific condition. Furthermore, based on Rademacher complexity, we prove that CLIP-ICM exhibits a lower generalization bound relative to CLIP, indicating the superior generalization performance of our method.
  • The experimental results verify that CLIP-ICM is very effective compared to the state-of-the-art and can effectively improve the performance of CLIP in dealing with zero-shot OOD problems. Meanwhile, we conduct a series of ablation studies to validate the effectiveness of our proposed CLIP-ICM.
2 Related Works
Vision-Language Pre-training.
Incorporating language with the visual modality has been a long-standing issue extensively researched (Frome et al., 2013; Socher et al., 2013; Norouzi et al., 2014; Elhoseiny et al., 2013). In recent years, multi-modal pre-training models led by CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021) have achieved significant success. The objective of this approach is to enable cross-modal learning by mapping language and visual modals into the same embedding space. Recently, researches have been focusing on improving the OOD performance when adapting CLIP to downstream tasks (Gao et al., 2023; Zhou et al., 2022a, b; Shu et al., 2023). (Gao et al., 2023) propose CLIP-Adapter to perform fine-tuning with feature adapters. (Zhou et al., 2022a, b) propose to learn dynamic prompts to improve the OOD performance of CLIP. (Shu et al., 2023) propose a fine-tuning method that adapts CLIP models to OOD situations. In our work, we analyze the reason for poor OOD generalization ability of previous methods. We propose CLIP-ICM to facilitate CLIP-related methods by learning invariant causal mechanisms.
Causal Representation Learning.
The latent factor hypothesis assumes that a set of latent factors generate the observational data (Bishop, 1998). A core task in causal representation learning is to provably identify these latent factors (Bengio et al., 2014; Schölkopf et al., 2021; Schoelkopf et al., 2012). Additionally, the issue of identification is commonly referred to as the Independent Component Analysis (ICA) problem (Hyvärinen & Oja, 2000), which aims to separate independent signals from those mixed during observation. While the problem of ICA has been proven to be achievable in the linear case (Hyvärinen & Oja, 2000), non-linear ICA has been shown to be impossible without inductive biases relying on auxiliary labels (Hyvärinen & Pajunen, 1999; Locatello et al., 2019; Hyvarinen & Morioka, 2017; Locatello et al., 2020), imposing sparsity (Schölkopf et al., 2021; Lachapelle et al., 2022), or restricting the function class (Zimmermann et al., 2022).
Importantly, many of these approaches are variations of the Variational Autoencoder framework (Kingma & Welling, 2013). Attention is also drawn to identification work on discriminative models (Zimmermann et al., 2022; Kügelgen et al., 2021; Roeder et al., 2021). Similar to our work, (Kügelgen et al., 2021) also uses data augmentations to obtain the invariant factors, and we provide a detailed comparison with them in Appendix H. Specifically, (Roeder et al., 2021) demonstrate that discriminative representation learning methods, including classification and contrastive learning (Oord et al., 2019; Radford et al., 2021), can identify latent factors with an affine transformation.(Daunhawer et al., 2023) also provide identifiability results for multimodal contrastive learning, demonstrating the recovery of shared factors in a more general multimodal setup beyond the multi-view settings previously studied.
While the studies mentioned above concentrate on scenarios with only observational data, an increasing number of studies are now focusing on cases where interventional data is available (Ahuja et al., 2023; Buchholz et al., 2023; Squires et al., 2023). We approach from a similar perspective and propose a method to identify latent factors through interventions in the data. Compared to the method they proposed, we introduce a simple and effective way to utilize intervention data to identify the invariant features of CLIP.
Out-of-distribution Generaliztion.
It is evident in many cases that the performance of deep learning methods is weakened when applied to different distributions (Beery et al., 2018; Taori et al., 2020). Studies have been conducted to address the OOD generalization issues (Arjovsky et al., 2020; Ahuja et al., 2020; Gulrajani & Lopez-Paz, 2020; Miller et al., 2021; Abbe et al., 2023; Chen et al., 2023). One important branch of work aims to find the invariant causal mechanisms across domains (Arjovsky et al., 2020; Peters et al., 2016; Suter et al., 2019; Ahuja et al., 2020, 2022a; Chen et al., 2023); the core of aforementioned works is inspired by the invariance principle from causality (Pearl, 2009; Pearl & Bareinboim, 2014; Schölkopf et al., 2021). (Robey et al., 2021; Wald et al., 2021; Li et al., 2024) also incorporate with SCM when analyzing the OOD problem, we provide a detailed comparison with them in Appendix H. Vision-language pre-trained models, such as CLIP (Radford et al., 2021), exhibit impressive performance across various domains. However, recent works emphasize that adapting CLIP with task-specific data comes at the cost of OOD generalization ability (Shu et al., 2023; Gao et al., 2023; Pham et al., 2023). In our investigation of the invariant causal mechanisms of CLIP, we theoretically demonstrate that our method exhibits superior OOD generalization ability.
3 Preliminaries
3.1 Contrastive Language-Image Pre-training
We begin by revisiting the framework for vision-language joint training, as introduced in (Radford et al., 2021). Let $\boldsymbol{x}\in\mathcal{X}$ be an arbitrary image, and $\boldsymbol{t}\in\mathcal{T}$ be a text sequence to describe this image, where $\mathcal{X}$ denotes the image space and $\mathcal{T}$ represents the text space. CLIP aims to train an image encoder $f_{I}$ and a text encoder $f_{T}$ to map $\boldsymbol{x}$ and $\boldsymbol{t}$ into a joint embedding space $f_{I}(\boldsymbol{x}),f_{T}(\boldsymbol{t})\in\mathbb{R}^{d}$. Given a mini-batch of $N$ image-text pairs $\{\boldsymbol{x}_{i},\boldsymbol{t}_{i}\}_{i=1}^{N}$, the objective for learning $f_{I}$ and $f_{T}$ is achieved by the following contrastive loss functions:
$\mathcal{L}_{I}$ $=-\frac{1}{N}\log\sum_{i=1}^{N}\frac{\exp(f_{I}(\boldsymbol{x}_{i})^{T}f_{T}(\boldsymbol{t}_{i})/\tau)}{\sum_{j=1}^{N}\exp(f_{I}(\boldsymbol{x}_{i})^{T}f_{T}(\boldsymbol{t}_{j})/\tau)},$ (1)
$\mathcal{L}_{T}$ $=-\frac{1}{N}\log\sum_{i=1}^{N}\frac{\exp(f_{T}(\boldsymbol{t}_{i})^{T}f_{I}(\boldsymbol{x}_{i})/\tau)}{\sum_{j=1}^{N}\exp(f_{T}(\boldsymbol{t}_{i})^{T}f_{I}(\boldsymbol{x}_{j})/\tau)},$ (2)
where $\tau$ is a temperature hyperparameter. Minimizing these two loss functions can maximize the similarity between the embeddings of the correct image pairs $\{f_{I}(\boldsymbol{x}_{i}),f_{T}(\boldsymbol{t}_{i})\}$, while minimizing the similarity between the wrong image pairs $\{f_{I}(\boldsymbol{x}_{i}),f_{T}(\boldsymbol{t}_{j})\},j\neq i$.
The CLIP model facilitates zero-shot predictions for unknown classification tasks with $C$ classes. For each class $c$, a corresponding text prompt $\boldsymbol{t}_{c}$ is employed, e.g., ”a photo of a [CLASS]” where [CLASS] represents the name of class $c$. The probability that an image sample $\boldsymbol{x}$ belongs to class $y$ is computed using the formula:
$p(y|\boldsymbol{x})=\frac{\exp(f_{I}(\boldsymbol{x})^{T}f_{T}(\boldsymbol{t}_{y}))}{\sum_{c=1}^{C}\exp(f_{I}(\boldsymbol{x})^{T}f_{T}(\boldsymbol{t}_{c}))}.$ (3)
Another commonly employed predictive approach, known as linear evaluation, aims to learn a classifier $W\in\mathbb{R}^{d\times C}$ with the parameters of the pre-trained model fixed. Using linear evaluation, the probability that a sample $\boldsymbol{x}$ belongs to class $y$ is calculated as follows:
$p(y|\boldsymbol{x};W)=\frac{\exp(W_{y,:}f_{I}(\boldsymbol{x}))}{\sum_{c=1}^{C}\exp(W_{c,:}f_{I}(\boldsymbol{x}))}.$ (4)
In this work, to assess the prediction performance of CLIP, we use $p_{\theta}(y|\boldsymbol{x})$ to denote the estimated probability by CLIP, where $\theta$ is the parameter.
3.2 Out-Of-Distribution Generalization
Next, we formalize the setting when CLIP is used across various domains. We consider supervised training data gathered from a set of environments $E_{tr}$: $\mathcal{D}=\{\mathcal{D}^{e}\}_{e\in E_{tr}}$, where $\mathcal{D}^{e}=\{\boldsymbol{x}_{i}^{e},y_{i}^{e}\}_{i=1}^{N^{e}}\sim P^{e}_{XY}$ represents the dataset from environment $e\in E_{tr}$. Here, $N^{e}$ is the number of instances in environment $e$, and $P^{e}_{XY}$ is the joint distribution of the sample and label in environment $e$, $\boldsymbol{x}_{i}^{e}$ is the image and $y_{i}^{e}$ is the corresponding class label. The goal of OOD generalization is to find a predictor $p_{\theta}(y|\boldsymbol{x})$, estimated by parameter $\theta$, that performs well across various unseen environments in $E_{all}$, where $E_{tr}\subset E_{all}$. Specifically, the objective is to minimize the OOD risk:
$R^{OOD}(\theta)=\max_{e\in E_{all}}R^{e}(\theta),$ (5)
where $R^{e}(\theta)=\mathbb{E}_{P^{e}_{XY}}[-\log p_{\theta}(y|\boldsymbol{x})]$ represents the risk of the predictor in environment $e$. The OOD typically encompasses two primary scenarios: domain shift, where the training and test data originate from different domains; and open class, where the test data introduces new classes not encountered during training. In the case of domain shift, the class label $y^{e}$ remains constant across all environments. Conversely, in open class scenarios, the distribution $P^{e}_{X}$ of the sample $\boldsymbol{x}$ remains fixed across all environments.
3.3 Background in Causality
The SCMs are used to describe the causal variables and their relationships. The definition of SCM is:
Definition 3.1.
(Structural Causal Model (Pearl, 2009))
1. A set $U$ of background or exogenous variables, representing factors outside the model, which nevertheless affect relationships within the model.
2. A set $V=\{V_{1},\dots,V_{n}\}$ of endogenous variables, assumed to be observable. Each of these variables is functionally dependent on some subset $PA_{i}$ of $U\cup V$. Here, $PA_{i}$ means the set of parent nodes of the endogenous variable $V_{i}$.
3. A set $F$ of functions $\{f_{1},\dots,f_{n}\}$ such that each $f_{i}$ determines the value $v_{i}$ of $V_{i}\in V$, $v_{i}=f_{i}(pa_{i},u)$. These functions are also known as causal mechanisms.
4. A joint probability distribution $P(u)$ over $U$.
The SCMs can be represented by Directed Acyclic Graphs (DAG) (Glymour et al., 2016). In the DAG, the arrow $X\rightarrow Y$ denotes that changes in the value of $X$ directly cause changes in $Y$. However, this relationship does not hold in the reverse order.
In statistical learning, we can only identify the association by $p(Y|X)$, indicating the correlation between variable $X$ and $Y$. However, in causal learning, our focus shifts to understanding the genuine causal mechanisms, specifically how the intervention of the variable $X$ influences another variable $Y$. This form of intervention is modeled by replacing the causal mechanisms $v_{i}=f_{i}(pa_{i},u)$ with a constant value, denoted by the $do$-operator (Bareinboim et al., 2022). In essence, while $p(Y|X)$ reflects statistical associations, $p(Y|do(X))$ signifies a causal relationship by explicitly considering the intervention in the causal structure.
4 Main Results
In this section, we first conduct a series of experiments using CLIP and observe various performances across different domains. Particularly, in certain domains, the classification results are notably poor. To investigate the prediction process of CLIP in various domains, we explicitly model the causal relationships between latent factors and observation data with SCM. Through analysis, we infer that the unsatisfactory results stem from the inherent lack of transferability of the causal mechanisms. Subsequently, we establish that the causal mechanism becomes transferable only when variant factors are observable, and the prediction depends solely on the invariant factors. Based on this insight, we propose to utilize interventional data to identify the latent factors. We propose CLIP-ICM, an algorithm to identify the invariant latent factors and conduct predictions based on them. Furthermore, we demonstrate that our method leads to a lower OOD generalization bound.
4.1 Motivation Experiment
We follow the standard zero-shot prediction protocol of CLIP (Radford et al., 2021) and report the prediction accuracy on different domains in the VLCS (Torralba & Efros, 2011) dataset and report the results in Table 1.
Table 1: Zero-shot and linear-probe prediction results in the VLCS dataset. Where C, L, S and V represent different domains.
Algorithm C L S V Avg
CLIP Zero-shot 99.9$\pm$0.0 70.1$\pm$0.0 73.5$\pm$0.0 86.1$\pm$0.0 82.4
CLIP Linear-probe 93.7$\pm$0.2 65.4$\pm$0.2 76.4$\pm$0.2 79.1$\pm$0.3 78.7
It is evident from the table that accuracy varies across different domains. Specifically, when performing prediction with both zero-shot and linear-probe manner, in the L domain, classification accuracy is notably lower compared to other domains. Consequently, according to Equation 5, the OOD risk of CLIP is high, indicating the unsatisfactory OOD generalization ability of CLIP.
4.2 Causal Analysis
(a) CLIP
(b) CLIP-ICM
Figure 2: The SCM of CLIP and CLIP-ICM in OOD scenarios. The hollow circles in (a) indicate that the corresponding variables are unobservable and the dashed lines indicate that the processes between two variables are unknown.
To further investigate the reason for the unsatisfactory OOD generalization ability of CLIP, we establish the SCM as in Figure 2(a). In the proposed SCM, $X$ is the image, and $Y$ is the text description. $Z_{inv}$ and $Z_{var}$ are latent factors that generate the image and label according to the latent factor hypothesis (Bishop, 1998). We also introduce an additional exogenous variable $E$ denotes the environment, i.e. the domain (Pearl & Bareinboim, 2014). $Z_{inv}$ is the set of latent factors that remain constant across various domains, and $Z_{var}$ is the set of latent factors that depend on environments. It is noteworthy that the variables $Z_{inv}$ and $Z_{var}$ are not observable. The edge $Z_{inv}\rightarrow X$ and the edge $Z_{var}\rightarrow X$ denote the generating process of the image from the latent factors. The edge $Z_{inv}\rightarrow Y$ and the edge $Z_{var}\rightarrow Y$ are the processes by which humans summarize features into abstract text. The edge $E\rightarrow Z_{var}$ denotes that $Z_{var}$ depends on the environments. The edge $X\rightarrow Y$ is the prediction process. The SCM is in the form of a selection diagram defined as follows:
Definition 4.1.
(Selection diagram): Let $<M,M^{*}>$ be a pair of SCMs relative to domain $<\Pi,\Pi^{*}>$, sharing a causal diagram $G$. $<M,M^{*}>$ is said to induce a selection diagram $D$ if $D$ is constructed as follows, a graphical illustration of the construction of the selection diagram is depicted in Figure 3:
1. Every edge in $G$ is also an edge in $D$
2. $D$ contains an extra edge $S_{i}\rightarrow V_{i}$ whenever there might exist a discrepancy $f_{i}\neq f_{i}^{*}$ or $P(U_{i})\neq P^{*}(U_{i})$ between $M$ and $M^{*}$.
The set $S$ are denoted as selection variables.
Figure 3: The construction of the selection diagram according to Definition 4.1. It can be observed that every edge in $G$ is also an edge in $D$. And $D$ contains an extra edge $E\rightarrow Z_{var}$ and $P(Z_{var})\neq P^{*}(Z_{var})$. The environment variable $E$ is the selection variable.
Consider the scenario where we estimate a well-performed predictor $p_{\theta}(y|\boldsymbol{x})$ in the domain $e_{1}$, which can be seen as an estimation of the distribution $P(Y|do(X),E=e_{1})$. We are interested in assessing under which circumstance, the predictor can still accurately predict $y$ from $x$ in an arbitrary domain $e_{2}$, i.e. accurately estimate the distribution $P(Y|do(X),E=e_{2})$. Achieving this criterion in prediction leads us to believe that the predictor demonstrates good generalization ability, and this criterion is often referred to as transportability.
Definition 4.2.
[…]
If the causal mechanism satisfies the following theorem, then it is considered transportable.
Theorem 4.3.
(sufficient and necessary conditions of transportability (Pearl & Bareinboim, 2014)) The average causal effect $P^{*}(Y|do(X))$ is transportable from $\Pi$ to $\Pi^{*}$ if either one of the following conditions holds:
1. $P^{*}(Y|do(X))$ is trivially transportable.
2. There exists a set of covariates $Z$ (possibly affected by $X$) such that $Z$ is $S$-admissible and for which $P^{*}(z|do(x))$ is transportable.
3. There exists a set of covariates, $W$ that satisfy $(X\perp\!\!\!\!\perp Y|W)_{\overline{X(W)}}$ and for which $P^{*}(w|do(x))$ is transportable.
The definition of trivially transportable is in Appendix A, and the definition of $S$-admissible is as follows:
Definition 4.4.
($S$-admissibility). A set $Z$ of variables satisfying $(Y\perp\!\!\!\!\perp S|Z,X)$ in $D_{\overline{X}}$ will be called S-admissible with respect to the causal effect of $X$ on $Y$. $D_{\overline{X}}$ denote deleting all arrows pointing to node $X$ in $D$.
If either one of the conditions of Theorem 4.3 is fulfilled, then it is possible to estimate $P(Y|do(X),E=e_{2})$, i.e. $P^{*}(Y|do(X))$ with our predictor. In the following, we are going to show, in Figure 2(a), unless the variables $Z_{var}$ and $Z_{inv}$ are observable, the causal relation $P^{*}(Y|do(X))$ is not transportable.
Theorem 4.5.
The formulation of the causal mechanisms $P^{*}(Y|do(X))$ in Figure 2(a) can be written as $\sum_{Z_{var}}\sum_{Z_{inv}}P(Y|do(X),Z_{var},Z_{inv})P(Z_{inv})P^{*}(Z_{var})$ (see Equation 15 in Appendix A for details). Where the $P(Y|do(X),Z_{inv})$ is the invariance causal mechanism that doesn’t depend on the environment $E$. However, without the observation of $Z_{inv}$, the causal mechanism $P^{*}(Y|do(X))$ in Figure 2(a) is not transportable from $\Pi$ to $\Pi^{*}$.
The detailed proof is provided in Appendix A. According to Theorem 4.5, $P(Y|do(X),Z_{inv})$ is the invariant causal mechanism, that can be estimated from the source domain $e_{1}$, while $P^{*}(Z_{var})$ is the distribution of $Z_{var}$ on target domain $e^{*}$. We also provide a detailed analysis of the invariant causal mechanism $P(Y|do(X),Z_{inv})$ in Appendix B. Theorem 4.5 state that estimating the target distribution required the observable of $Z_{inv}$. In the next subsection, we provide a way to provably find (identify) these factors.
4.3 Methodology
According to Theorem 4.5, in order to capture the causal mechanism $P(Y|do(X),E=e)$, access to $Z_{var}$ is required. This challenge mirrors the well-established problem of definitively identifying the latent factors responsible for generating the data (Schölkopf et al., 2021; Hyvärinen & Oja, 2000). In the following subsection, we demonstrate that by leveraging interventional data, it is possible to identify these latent factors up to permutation and scaling. Building upon this insight, we propose CLIP-ICM to idenfify them.
We begin by considering the latent factor $Z\in\mathbb{R}^{d}$, sampled from the distribution $P_{Z}$ supported on $\mathcal{Z}$. The observed data $x\in\mathbb{R}^{n}$ is generated from $Z$ through an injective mapping $g:\mathbb{R}^{d}\rightarrow\mathbb{R}^{n}$. The distribution of observed data is $P_{X}$ supported on $\mathcal{X}=g(\mathcal{Z})$. Our goal is to learn an encoder function $f_{\theta}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}$, where $\theta$ represents its parameters. Ideally, the encoder function should invert the data-generating process, satisfying $f_{\theta^{\prime}}(\boldsymbol{x})=Z$. However, due to the absence of additional knowledge about the latent distribution $P_{Z}$, we can only prove that, under a specific constraint, the encoder function’s output is a linear transformation of the true latent $Z$. Specifically, $f_{\theta}(\boldsymbol{x})=AZ$, where $A\in\mathbb{R}^{d\times d}$ is an invertible matrix.
Proposition 4.6.
For an injective data generating process $\boldsymbol{x}=g(Z)$, suppose there exists an ideal encoder $f_{\theta^{\prime}}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}$, such that $f_{\theta^{\prime}}(\boldsymbol{x})=Z$. If for all environment $E_{all}$, the encoder is trained optimal, and there exists a predictor for each environment, such that $p_{\theta}(y^{e}|\boldsymbol{x}^{e};W_{\phi_{e}})=p_{\theta^{\prime}}(y^{e}|\boldsymbol{x}^{e};W^{e}_{\phi^{\prime}_{e}}),\forall e\in E_{all}$. Here, $W_{\phi_{e}}$ denote the classifier trained based on $\theta$ on environment $e$, where $\phi_{e}$ denotes the parameters of this classifier. $W^{e}_{\phi^{\prime}_{e}}$ denote the classifier trained based on $\theta^{\prime}$ on environment $e$, where $\phi_{e}^{\prime}$ denote the parameters of this classifier. Then the encoder $f_{\theta}(\boldsymbol{x})$ identifies the latent factor $Z$ up to an invertible linear transformation, i.e. $f_{\theta}(\boldsymbol{x})=AZ$.
Figure 4: A graphical illustration of Proposition 4.6. In the figure, suppose that the image $\boldsymbol{x}$ is generated from the latent space $Z=[Z_{inv},Z_{var,1},Z_{var,2}]$ with $g(Z)$. Proposition 4.6 implies that the output of the CLIP encoder $f_{\theta}(\boldsymbol{x})=\hat{Z}=[\hat{Z}_{1},\hat{Z}_{2},\hat{Z}_{3}]$ is a invertible affine transformation of $Z$.
The proof is illustrated in Appendix C. And Figure 4 provides a graphical explanation. This conclusion is a commonly recognized result from a series of works (Roeder et al., 2021; Ahuja et al., 2022b; Hyvarinen & Morioka, 2016; Zimmermann et al., 2022), and Proposition 4.6 provides a reproof of it in our scenario. CLIP can be considered satisfying Proposition 4.6 due to its impressive performance as a powerful pre-trained model. It was trained on a dataset of 400 million (image, text) pairs collected from the Internet and has demonstrated success in linear evaluation on downstream tasks (Radford et al., 2021). However, the linear identifiability of latent factors alone is insufficient to ensure the transportability of causal effects. This limitation arises from the fact that the mixing matrix $A$ is arbitrary, hindering the identification of domain-related factors $Z_{var}$.
Note that the aforementioned analyses are all based on observational data. The causal mechanism seeks to answer questions related to interventions, such as predicting the model’s response when changing a real-world image to a sketched painting. This involves intervening on factors like background color and texture. The operation, known as a $do$-intervention, alters latent factors, and the resulting distribution of $Z$ differs from the observational distribution $P_{Z}$. Formally, a $do$-intervention sets the value of a specific latent factor to a fixed value $do(z_{i})=z^{\dagger}$, while the distribution of other latent factors is $P^{do(z_{i})}_{Z}$. The distribution of the interventional data is $P^{do(z_{i})}_{X}$ supported on $\mathcal{X}^{do(z_{i})}$.
Next, we demonstrate that by intervening on a specific latent factor, denoted as $z_{i}$, such that for all $\boldsymbol{x}\in\mathcal{X}^{do(z_{i})}$, the $k$-th output of the encoder remains a fixed value, the encoder can identify the corresponding intervened latent factor up to shift and scaling.
Theorem 4.7.
Consider observational data generated from an injective mapping $\boldsymbol{x}=g(Z)$, where the latent factors $Z$ follow a distribution $P_{Z}$ supported on $\mathcal{Z}$. If a $do$-intervention is applied to the latent factor $z_{i}$, with the distribution of other latent factors denoted as $P^{do(z_{i})}_{Z}$, the resulting distribution of the interventional data is denoted as $P^{do(z_{i})}_{X}$. Suppose that an encoder $f_{\theta}$ satisfies Proposition 4.6, and the output of the $k^{th}$ component of the encoder is required to take a fixed value $f_{\theta}(\boldsymbol{x})[k]=z^{\dagger}$ for all $x\in\mathcal{X}^{do(z_{i})}$. Then, the encoder identifies the intervened latent $z_{i}$ up to a shift and scaling, i.e., $f_{\theta}(\boldsymbol{x})[k]=a_{i}z_{i}+b_{i}$, where $a_{i}\in\mathbb{R}$ and $b_{i}\in\mathbb{R}$.
The proof is detailed in Appendix D. Theorem 4.7 presents an alternative approach for accessing $Z_{var}$ and $Z_{inv}$. When we intervene on all latent factors, the encoder identifies these factors up to permutation and scaling. This transformation alters the SCM from Figure 2(a) to Figure 2(b). As per Definition 4.4 and Theorem 4.3, the invariant factors $Z_{inv}$ is identified, this implies that the causal mechanism $P(Y|do(X),Z_{inv})$ in Figure 2(b) is transportable. In other words, the encoder captures the invariant causal mechanism.
Theorem 4.7 provides a condition under which $Z_{inv}$ can be identified. Subsequently, leveraging the condition outlined in Theorem 4.7, we introduce a straightforward and efficient algorithm named CLIP-ICM. This algorithm functions as a post-processing step for the output of the CLIP encoders and is designed to meet the requirements of the condition specified in Theorem 4.7 We perform a set of interventions by utilizing data augmentation techniques, including color distortion, grayscale conversion, and Gaussian blur, on the raw image. Each type of data augmentation is regarded as an element of $Z_{var}$. For an image sample with different augmentations, it can be regarded that only the value of $Z_{var}$ is changed while the $Z_{inv}$ remains a fixed value $z^{\dagger}$. Therefore, according to Theorem 4.7, we propose to estimate a linear transformation $C_{\xi}\in\mathbb{R}^{D_{inv}\times D}$ to produce the $Z_{inv}$, where a total of $K$ augmentations are employed. Empirically, $D_{inv}$ is the dimension of the estimated $Z_{inv}$ space, which is a hyper-parameter. We provide a detailed ablation study on this hyperparameter in Appendix I. The $C_{\xi}$ should exhibit two important characteristics: (1) the transformed representation $C_{\xi}f_{\theta}(\boldsymbol{x})$ should contain as much information about the original representation $f_{\theta}(\boldsymbol{x})$ as possible. (2) $C_{\xi}f_{\theta}(\boldsymbol{x})$ should be invariant to any intervention on $Z_{var}$. Let $\hat{Z}=f_{\theta}(\boldsymbol{x})$, the objective for $C_{\xi}$ is:
$\begin{array}[]{l}\mathop{\min}\limits_{{C_{\xi}}}\|{\hat{Z}-{{\hat{Z}}^{T}}C_{\xi}^{T}{C_{\xi}}}\|_{2}^{2}+\lambda\sum\limits_{j\in K}{\|{{C_{\xi}}(\hat{Z}-{f_{\theta}}({\alpha_{j}}(\boldsymbol{x})))}\|_{2}^{2}}\\ s.t.\quad\quad rank({{\hat{Z}}^{T}}C_{\xi}^{T}{C_{\xi}})\leq D_{inv},\end{array}$
where $\lambda$ is a hyperparameter, $j$ is the type of data augmentation, $\alpha_{j}$ is the $j$-th data augmentation and $f_{\theta}$ is the feature extractor.
Figure 5: Graphical illustration of Equation 6. Minimizing Equation 6 is equivalent to finding the principle components of $\hat{Z}$ which is invariant to data augmentations $\alpha_{j}$, i.e. finding the normal vector of the $Z_{var}$ surface.
Equation 6 can be viewed as a low-rank approximation problem (Golub et al., 1987) if we view the goal of finding a mapping from $\hat{Z}$ to $Z_{inv}$ through the lens of principle component analysis (Krogh & Hertz, 1991). We consider the true latent space of $Z$ as a composite of two orthogonal subspaces, i.e. $Z=Z_{inv}\times Z_{var}$. According to Proposition 4.6, the learned representation space $\hat{Z}=AZ$ is an invertible linear transformation of the true latent space. Therefore, to find a mapping from $\hat{Z}$ to $Z_{inv}$, is equivalent to finding the principle component of $\hat{Z}$ which is orthogonal to any intervention on $Z_{var}$, i.e. invariant to data augmentation. This is the idea of the second term in Equation 6. Empirically, we wish the invariant feature to contain as much information about the information of image as possible, and that is the idea of the first term in Equation 6.
To perform the invariant prediction, the invariant zero-shot prediction is obtained as follows:
$p(y|do(\boldsymbol{x}),Z_{inv})=\frac{\exp((C_{\xi}f_{I}(\boldsymbol{x}))^{T}(C_{\xi}f_{T}(\boldsymbol{t}_{y})))}{\sum_{c=1}^{C}\exp((C_{\xi}f_{I}(\boldsymbol{x}))^{T}(C_{\xi}f_{T}(\boldsymbol{t}_{c})))},$
where the predictions are solely based on the similarities derived from $Z_{inv}$. For the linear probe evaluation, the $C_{\xi}$ only applied to the output of the image encoder, and we train a classifier accordingly.
4.4 Generalization Analysis
In this subsection, we prove that the OOD generalization ability of our method is better than CLIP based on Rademacher complexity (Mohri et al., 2018), i.e., CLIP-ICM has a lower generalization bound compared to CLIP.
Let $Z=[z_{var,1},\dots,z_{var,m},z_{inv,1},\dots,z_{inv,n}]\in\mathbb{R}^{d}$ be the latent factors that generate observation data $\boldsymbol{x}=g(Z)$. According to Proposition 4.6, the output of the CLIP image encoder is $f_{\theta}(\boldsymbol{x})=AZ$, where $A$ is an invertible matrix. Assuming that the classification depends only on the invariant part, i.e. $z_{inv}$, we aim to learn a classifier $W_{Z}$ based on the latent factor matrix $Z$. Under the above assumption, $W_{Z}$ is a sparse matrix, while for the entangled latent factor matrix $\hat{Z}=AZ$, the corresponding $W_{\hat{Z}}$ is dense.
For a hypothesis space $H$ and a data distribution $\mathcal{D}$, Rademacher complexity (Mohri et al., 2018) is a powerful metric to measure the OOD generalization ability of functions in $H$. Denote by $R^{OOD}(h)$ the OOD generalization error of hypothesis $h\in H$, our main conclusion could be summarized as follows:
Theorem 4.8.
For any data distribution $\mathcal{D}$, finite hypothesis spaces $H_{Z},H_{\hat{Z}}\subset\mathbb{R}^{m}$, the generalization error of hypothesis $W_{Z}\in H_{Z}$ and hypothesis $W_{\hat{Z}}\in H_{\hat{Z}}$ satisfies the following inequality:
$R^{OOD}(H_{Z},\mathcal{D})\leq R^{OOD}(H_{\hat{Z}},\mathcal{D}).$ (8)
Theorem 4.8 indicates that under identical confidence level, the OOD generalization error of $Z$ is lower than $\hat{Z}$, which ensures a better OOD generalization ability of CLIP-ICM. The proof is illustrated in Appendix E.
5 Experiment
Table 2: Accuracy on VLCS dataset. C,L,S and V represents different domains.
Algorithm Backbone C L S V Avg
ERM R50 97.7$\pm$0.4 64.3$\pm$0.9 73.4$\pm$0.5 74.6$\pm$1.3 77.5
IRM (Arjovsky et al., 2020) R50 98.6$\pm$0.1 64.9$\pm$0.9 73.4$\pm$0.6 77.3$\pm$0.9 78.5
GroupDRO (Sagawa* et al., 2019) R50 97.3$\pm$0.3 63.4$\pm$0.9 69.5$\pm$0.8 76.7$\pm$0.7 76.7
Mixup (Yan et al., 2020) R50 98.3$\pm$0.6 64.8$\pm$1.0 72.1$\pm$0.5 74.3$\pm$0.8 77.4
MMD (Li et al., 2018) R50 97.7$\pm$0.1 64.0$\pm$1.1 72.8$\pm$0.2 75.3$\pm$3.3 77.5
DANN (Ganin et al., 2016) R50 99.0$\pm$0.3 65.1$\pm$1.4 73.1$\pm$0.3 77.2$\pm$0.6 78.6
ARM (Zhang et al., 2021) R50 98.7$\pm$0.2 63.6$\pm$0.7 71.3$\pm$1.2 76.7$\pm$0.6 77.6
MBDG (Robey et al., 2021) R50 98.3 68.1 68.8 76.3 77.9
CLIP Linear-probe R50 98.1$\pm$0.7 63.8$\pm$0.8 79.8$\pm$0.7 84.5$\pm$1.0 81.5
CLIP-ICM Linear-probe R50 99.2$\pm$0.1 69.9$\pm$0.3 80.5$\pm$0.3 89.5$\pm$0.9 84.8
CLIPOOD (Shu et al., 2023) ViT-B 97.5$\pm$0.6 68.3$\pm$0.5 83.9$\pm$0.9 88.7$\pm$1.0 84.6
CLIP Linear-probe ViT-B 93.7$\pm$0.2 65.4$\pm$0.2 76.4$\pm$0.2 79.1$\pm$0.3 78.7
CLIP-ICM Linear-probe ViT-B 98.8$\pm$0.3 72.7$\pm$0.8 86.4$\pm$0.9 87.9$\pm$0.9 86.5
CLIP Zero-shot R50 99.2$\pm$0.0 69.5$\pm$0.0 69.8$\pm$0.0 84.9$\pm$0.0 80.9
CLIP-ICM Zero-shot R50 99.4$\pm$0.4 71.8$\pm$0.8 71.7$\pm$0.8 85.3$\pm$1.0 82.1
CLIP Zero-shot ViT-B 99.9$\pm$0.0 70.1$\pm$0.0 73.5$\pm$0.0 86.1$\pm$0.0 82.4
CLIP-ICM Zero-shot ViT-B 100.0$\pm$0.0 74.7$\pm$0.4 74.5$\pm$0.4 87.1$\pm$0.7 84.1
Table 3: Accuracy on PACS dataset. A,C,P and S represents different domains.
Algorithm Backbone P A C S Avg
ERM R50 97.2$\pm$0.3 84.7$\pm$0.4 80.8$\pm$0.6 79.3$\pm$1.0 85.5
IRM (Arjovsky et al., 2020) R50 96.7$\pm$0.6 84.8$\pm$1.3 76.4$\pm$1.1 76.1$\pm$1.0 83.5
GroupDRO (Sagawa* et al., 2019) R50 96.7$\pm$0.3 83.5$\pm$0.9 79.1$\pm$0.6 78.3$\pm$2.0 84.4
Mixup (Yan et al., 2020) R50 97.6$\pm$0.1 86.1$\pm$0.5 78.9$\pm$0.8 75.8$\pm$1.8 84.6
MMD (Li et al., 2018) R50 96.6$\pm$0.2 86.1$\pm$1.4 79.4$\pm$0.9 76.5$\pm$0.5 84.6
DANN (Ganin et al., 2016) R50 97.3$\pm$0.4 86.4$\pm$0.8 77.4$\pm$0.8 73.5$\pm$2.3 83.6
ARM (Zhang et al., 2021) R50 97.4$\pm$0.3 86.8$\pm$0.6 76.8$\pm$0.5 79.3$\pm$1.2 85.1
MBDG (Robey et al., 2021) R50 97.0 80.6 79.3 85.2 85.6
CLIP Linear-probe R50 99.2$\pm$0.7 91.7$\pm$0.8 92.5$\pm$0.7 82.3$\pm$0.2 91.4
CLIP-ICM Linear-probe R50 99.4$\pm$0.5 92.8$\pm$0.7 93.9$\pm$0.4 83.8$\pm$0.5 92.5
CLIP Linear-probe ViT-B 97.3$\pm$0.7 98.4$\pm$0.1 99.5$\pm$0.4 90.4$\pm$0.3 96.4
CLIP-ICM Linear-probe ViT-B 98.8$\pm$0.2 99.4$\pm$0.3 99.8$\pm$0.8 92.1$\pm$0.5 97.5
CLIP Zero-shot R50 99.3$\pm$0.0 91.0$\pm$0.0 93.1$\pm$0.0 80.5$\pm$0.0 91.0
CLIP-ICM Zero-shot R50 99.4$\pm$0.4 93.6$\pm$0.2 99.3$\pm$0.5 83.1$\pm$0.6 93.8
CLIP Zero-shot ViT-B 100.0$\pm$0.0 97.4$\pm$0.0 99.2$\pm$0.0 88.1$\pm$0.0 96.1
CLIP-ICM Zero-shot ViT-B 98.7$\pm$0.5 99.1$\pm$0.4 99.9$\pm$0.1 91.7$\pm$0.4 97.3
Table 4: Accuracy on Office-Home dataset. A,C,P and R represents different domains.
Algorithm Backbone A C P R Avg
ERM R50 61.3$\pm$0.7 52.4$\pm$0.3 75.8$\pm$0.1 76.6$\pm$0.3 66.5
IRM (Arjovsky et al., 2020) R50 58.9$\pm$2.3 52.2$\pm$1.6 72.1$\pm$2.9 74.0$\pm$2.5 64.3
GroupDRO (Sagawa* et al., 2019) R50 60.4$\pm$0.7 52.7$\pm$1.0 75.0$\pm$0.7 76.0$\pm$0.7 66.0
Mixup (Yan et al., 2020) R50 62.4$\pm$0.8 54.8$\pm$0.6 76.9$\pm$0.3 78.3$\pm$0.2 68.1
MMD (Li et al., 2018) R50 60.4$\pm$0.2 53.3$\pm$0.3 74.3$\pm$0.1 77.4$\pm$0.6 66.3
DANN (Ganin et al., 2016) R50 59.9$\pm$1.3 53.0$\pm$0.3 73.6$\pm$0.7 76.9$\pm$0.5 65.9
ARM (Zhang et al., 2021) R50 58.9$\pm$0.8 51.0$\pm$0.5 74.1$\pm$0.1 75.2$\pm$0.3 64.8
SIG (Li et al., 2024) R50 76.4 63.9 85.4 85.8 77.8
CLIP Linear-probe R50 68.0$\pm$0.8 46.3$\pm$0.3 80.4$\pm$0.9 81.9$\pm$0.7 69.1
CLIP-ICM Linear-probe R50 78.3$\pm$0.3 56.4$\pm$0.8 88.6$\pm$0.8 87.7$\pm$0.7 77.8
CLIP Linear-probe ViT-B 78.9$\pm$0.9 69.3$\pm$0.9 90.3$\pm$0.3 89.0$\pm$0.2 81.9
CLIP-ICM Linear-probe ViT-B 84.3$\pm$0.5 71.4$\pm$0.2 92.5$\pm$0.2 90.2$\pm$0.8 84.6
CLIP Zero-shot R50 71.3$\pm$0.0 50.4$\pm$0.0 81.7$\pm$0.0 82.6$\pm$0.0 71.5
CLIP-ICM Zero-shot R50 72.6$\pm$0.4 55.0$\pm$0.9 83.2$\pm$0.3 83.7$\pm$0.8 73.6
CLIP Zero-shot ViT-B 83.3$\pm$0.0 65.3$\pm$0.0 89.0$\pm$0.0 89.3$\pm$0.0 81.7
CLIP-ICM Zero-shot ViT-B 83.7$\pm$0.7 67.4$\pm$0.6 90.8$\pm$0.9 90.1$\pm$0.6 82.6
Table 5: Accuracy on TerraIncognita dataset. L100,L38,L43 and L46 represents different domains.
Algorithm Backbone L100 L38 L43 L46 Avg
ERM R50 49.8$\pm$4.4 42.1$\pm$1.4 56.9$\pm$1.8 35.7$\pm$3.9 46.1
IRM (Arjovsky et al., 2020) R50 54.6$\pm$1.3 39.8$\pm$1.9 56.2$\pm$1.8 39.6$\pm$0.8 47.6
GroupDRO (Sagawa* et al., 2019) R50 41.2$\pm$0.7 38.6$\pm$2.1 56.7$\pm$0.9 36.4$\pm$2.1 43.2
Mixup (Yan et al., 2020) R50 59.6$\pm$2.0 42.2$\pm$1.4 55.9$\pm$0.8 33.9$\pm$1.4 47.9
MMD (Li et al., 2018) R50 41.9$\pm$3.0 34.8$\pm$1.0 57.0$\pm$1.9 35.2$\pm$1.8 42.2
DANN (Ganin et al., 2016) R50 51.1$\pm$3.5 40.6$\pm$0.6 57.4$\pm$0.5 37.7$\pm$1.8 46.7
ARM (Zhang et al., 2021) R50 49.3$\pm$0.7 38.3$\pm$2.4 55.8$\pm$0.8 38.7$\pm$1.3 45.5
CLIP Linear-probe R50 52.0$\pm$1.0 34.4$\pm$0.6 56.1$\pm$0.9 32.8$\pm$0.2 44.1
CLIP-ICM Linear-probe R50 64.1$\pm$0.4 46.5$\pm$0.7 61.1$\pm$0.6 42.8$\pm$0.4 53.6
CLIPOOD (Shu et al., 2023) ViT-B 74.6$\pm$0.3 57.3$\pm$0.5 59.1$\pm$0.9 47.7$\pm$0.9 59.7
CLIP Linear-probe ViT-B 73.6$\pm$0.5 58.3$\pm$0.7 61.0$\pm$0.1 47.8$\pm$1.0 60.2
CLIP-ICM Linear-probe ViT-B 78.7$\pm$0.5 60.2$\pm$0.9 66.0$\pm$0.8 52.4$\pm$0.1 64.3
CLIP Zero-shot R50 7.7$\pm$0.0 14.8$\pm$0.0 32.4$\pm$0.0 20.9$\pm$0.0 19.0
CLIP-ICM Zero-shot R50 38.8$\pm$0.4 33.6$\pm$0.6 33.6$\pm$0.1 30.1$\pm$0.7 34.0
CLIP Zero-shot ViT-B 52.0$\pm$0.0 20.4$\pm$0.0 32.8$\pm$0.0 31.6$\pm$0.0 34.2
CLIP-ICM Zero-shot ViT-B 59.3$\pm$0.4 46.2$\pm$0.4 52.4$\pm$0.7 41.6$\pm$0.8 49.9
Table 6: Accuracy on DomainNet dataset.
Algorithm Backbone CLIPART INFOGRAPH PAINTING QUICKDRAW REAL SKETCH Avg
ERM R50 58.1$\pm$0.3 18.8$\pm$0.3 46.7$\pm$0.3 12.2$\pm$0.4 59.6$\pm$0.1 49.8$\pm$0.4 40.9
IRM (Arjovsky et al., 2020) R50 48.5$\pm$2.8 15.0$\pm$1.5 38.3$\pm$4.3 10.9$\pm$0.5 48.2$\pm$5.2 42.3$\pm$3.1 33.9
GroupDRO (Sagawa* et al., 2019) R50 47.2$\pm$0.5 17.5$\pm$0.4 33.8$\pm$0.5 9.3$\pm$0.3 51.6$\pm$0.4 40.1$\pm$0.6 33.3
Mixup (Yan et al., 2020) R50 55.7$\pm$0.3 18.5$\pm$0.5 44.3$\pm$0.5 12.5$\pm$0.4 55.8$\pm$0.3 48.2$\pm$0.5 39.2
MMD (Li et al., 2018) R50 32.1$\pm$13.3 11.0$\pm$4.6 26.8$\pm$11.3 8.7$\pm$2.1 32.7$\pm$13.8 28.9$\pm$11.9 23.4
DANN (Ganin et al., 2016) R50 53.1$\pm$0.2 18.3$\pm$0.1 44.2$\pm$0.7 11.8$\pm$0.1 55.5$\pm$0.4 46.8$\pm$0.6 38.3
ARM (Zhang et al., 2021) R50 49.7$\pm$0.3 16.3$\pm$0.5 40.9$\pm$1.1 9.4$\pm$0.1 53.4$\pm$0.4 43.5$\pm$0.4 35.5
CLIP Linear-probe R50 59.2$\pm$0.1 38.6$\pm$0.9 54.1$\pm$0.7 8.2$\pm$0.7 56.4$\pm$0.7 39.6$\pm$0.7 42.6
CLIP-ICM Linear-probe R50 63.1$\pm$0.1 50.1$\pm$0.2 66.2$\pm$1.0 19.7$\pm$0.3 83.8$\pm$0.4 63.2$\pm$0.3 57.8
CLIPOOD (Shu et al., 2023) ViT-B 77.6 54.7 72.5 20.7 85.7 69.9 63.5
CLIP Linear-probe ViT-B 73.9$\pm$0.5 40.2$\pm$0.5 62.1$\pm$0.2 15.1$\pm$0.3 76.9$\pm$0.1 62.0$\pm$0.7 55.0
CLIP-ICM Linear-probe ViT-B 78.6$\pm$0.2 55.6$\pm$0.2 72.9$\pm$0.6 22.1$\pm$0.7 86.1$\pm$0.3 68.5$\pm$0.3 64.0
CLIP Zero-shot R50 53.1$\pm$0.0 39.2$\pm$0.0 52.9$\pm$0.0 5.7$\pm$0.0 76.7$\pm$0.0 48.0$\pm$0.0 45.9
CLIP-ICM Zero-shot R50 58.0$\pm$0.1 42.3$\pm$0.5 54.4$\pm$0.2 12.8$\pm$0.7 78.2$\pm$0.1 49.6$\pm$0.2 49.2
CLIP Zero-shot ViT-B 70.2$\pm$0.0 46.6$\pm$0.0 65.0$\pm$0.0 13.7$\pm$0.0 82.9$\pm$0.0 62.7$\pm$0.0 56.8
CLIP-ICM Zero-shot ViT-B 77.1$\pm$0.5 51.8$\pm$0.5 67.9$\pm$1.0 16.7$\pm$0.3 82.9$\pm$0.8 66.9$\pm$0.2 60.5
In this section, we first present the implementation details of our proposed CLIP-ICM, and present the classification results in the OOD datasets. Then, we conduct ablation studies of our proposed method. Experimental results demonstrate the effectiveness of our proposed method.
5.1 Implementation Details.
We employ 7 data augmentation techniques, i.e. ColorJitter, GrayScale, GaussianBlur, RandomInvert, RandomPosterize, RandomSolarize, and RandomEqualize for producing interventional data. The matrix $C_{\xi}$ is obtained by solving a low-rank approximation problem through singular value decomposition. For the linear-probe evaluation, the classifier undergoes training for a maximum of 1000 epochs with an $l_{1}$ penalty of $1e^{-4}$ implementing with the built-in function of Scikit-learn (Kramer, 2016). It is noteworthy that we learn $C_{\xi}$ in a totally unsupervised manner without the inclusion of label information. Additionally, during the estimating of $C_{\xi}$, there are no images from the target domain included.
5.2 Classification in OOD datasets
Benchmarks and Evaluation Protocols.
To validate the effectiveness of the proposed CLIP-ICM in the OOD scenarios, we conduct experiments on the well-known benchmark DomainBed (Gulrajani & Lopez-Paz, 2020). Specifically, we select 5 datasets from DomainBed, namely VLCS (Fang et al., 2013), PCAS (Li et al., 2017), Office-Home(Venkateswara et al., 2017), TerraIncognita (Beery et al., 2018), and DomainNet (Peng et al., 2019). Each of these datasets contains the same categories across different domains. We utilize two types of evaluation protocols for our proposed CLIP-ICM, namely the zero-shot prediction and linear-probe prediction following (Radford et al., 2021). We follow the leave-one-out evaluation protocol, where one domain is chosen as the test domain and the other domains are chosen as the training domains.
Results.
The classification results of CLIP-ICM on VLCS, PCAS, Office-Home, TerraIncognita, and Domainnet are illustrated in Table 2, Table 3, Table 4, Table 5 and Table 6 respectively. It is clear from the tables that the CLIP-ICM significantly improves the performance of CLIP in both zero-shot and linear-probe manner. Specifically, in the linear probe of the VLCS dataset, CLIP-ICM shows a notable improvement of 10.0% in the S domain. Similarly, in the zero-shot setting of the TerraIncognita dataset, CLIP-ICM achieves an impressive improvement of 31.1% in the L100 domain. Moreover, the accuracy of CLIP-ICM is consistently improved across all domains, surpassing other state-of-the-art domain generalization methods.
Figure 6: The linear-probe accuracy of CLIP-ICM on the LabelMe domain of VLCS dataset
Figure 7: The zero-shot of CLIP-ICM accuracies on the VLCS dataset with different $\lambda$.
5.3 Ablation Study
5.4 The impact of different data augmentation
We investigate how CLIP-ICM performs with varying data augmentation strategies on the VLCS dataset. The experiments include different combinations of augmentation techniques and the zero-shot prediction accuracy on the LabelMe domain is illustrated in Figure 6. The diagonal of Figure 6 indicates scenarios where only one augmentation is applied, while other cells represent combinations of data augmentation techniques. It is noteworthy that the use of data augmentation significantly improves zero-shot performance, with certain combinations showing even more promising results.
5.5 The impact of hyperparameter $\lambda$.
We conduct experiments on the VLCS dataset and visualize the results in Figure 7. From the results, we can observe that $\lambda$ has a significant impact on the prediction results. This is because that higher value of $\lambda$ indicates the second term of Equation 6 is more important and has a higher impact. These results illustrating the effectiveness of our proposed method.
6 Conclusion
In our work, we observe various performances of CLIP across different domains, noticing both well and unsatisfactory performance across domains. Through causal analysis, we demonstrate the inadequacy of the current CLIP model in capturing invariant causal mechanisms. Furthermore, we prove that the invariant causal mechanism is identifiable only when variant and invariant latent factors are identifiable. Consequently, we propose the CLIP-ICM methodology based on theoretical support derived from identifying latent factors with the aid of intervention data. Leveraging data augmentation, we incorporate a linear transformation onto the CLIP pre-trained model to acquire invariant latent factors. Theoretical evidence shows that our approach outperforms CLIP in OOD generalization. Experimental results demonstrate the superior results of CLIP-ICM on the Domainbed Benchmark in comparison to existing methods.
Impact Statements
We believe that this work is a step towards general artificial intelligence. This technique introduces the invariant causal mechanism into vision-language models, potentially revealing the physical world’s essence and promoting interpretable machine learning with significant societal implications.
Appendix
This supplementary material provides detailed proofs for the theorem and proposition mentioned in the main text. Furthermore, additional experimental results and the implementation detailed are provided.
Appendix A Proof for Theorem 4.5
Before we proceed to the proof of Theorem 4.5, we first provide some basic definitions of causal inference.
Definition A.1.
($d$-separation): A set $S$ of nodes is said to block a path $p$ if either
1. $p$ contains at least one arrow-emitting node that is in $S$,
2. $p$ contains at least one collision node that is outside $S$ and has no descendant in $S$.
If $S$ blocks all paths from set $X$ to set $Y$, it is said to $d$-separate $X$ and $Y$. $X$ and $Y$ are independent given $S$, written $X\perp\!\!\!\!\perp Y|S$.
The $d$-separation reflects the conditional independencies that hold the distribution $P$ that is compatible with the DAG.
Definition A.2.
(Rules of do-calculus) Let $X,Y,Z,W$ be arbitrary disjoint sets of nodes in a causal DAG $G$. We denote by $G_{\overline{X}}$ the graph obtained by deleting from $G$ all arrows pointing to nodes in $X$. Likewise, we denote by $G_{\underline{X}}$ the graph obtained by deleting from $G$ all arrows emerging from nodes in $X$. The notation $G_{\overline{X}\underline{Z}}$ represents the deletion of both incoming and outgoing arrows.
1. Insertion/deletion of observations.
$P(Y|do(X),Z,W)=P(Y|do(X),W)\quad\text{if }(Y\perp\!\!\!\!\perp Z|X,W)_{G_{\overline{X}}}.$
2. Action/observation exchange.
$P(Y|do(X),do(Z),W)=P(Y|do(X),W)\quad\text{if }(Y\perp\!\!\!\!\perp Z|X,W)_{G_{\overline{X}\underline{Z}}}.$
3. Insertion/deletion of actions.
$P(Y|do(X),do(Z),W)=P(Y|do(X),W)\quad\text{if }(Y\perp\!\!\!\!\perp Z|X,W)_{G_{\overline{XZ(W)}}}.$
where $Z(W)$ is the set of $Z$-nodes that are not ancestors of any $W$-node in $G_{\overline{X}}$.
Definition A.3.
(Identifiability) A causal query $Q(M)=P(Y|do(X))$ is identifiable, given a set of assumptions $A$, if for any two (fully specified) models, $M_{1}$ and $M_{2}$ that satisfy $A$, we have:
$P(M_{1})=P(M_{2})\Rightarrow Q(M_{1})=Q(M2)$
Definition A.4.
(Trivial transportability). A causal relation $R$ is said to be trivially transportable from $\Pi$ to $\Pi^{*}$, if $R(\Pi^{*})$ is identifiable from $(G^{*},P^{*})$.
$R$ is trivially transportable if we can directly estimate $R(\Pi^{*})$ from observational data of $\Pi^{*}$, unaided by the causal information from $\Pi$. The following state the sufficient and necessary conditions of transportability of average causal effect $P^{*}(Y|do(X))$.
Theorem A.5.
(sufficient and necessary conditions of transportability (Pearl & Bareinboim, 2014)) The average causal effect $P^{*}(Y|do(X))$ is transportable from $\Pi$ to $\Pi^{*}$ if either one of the following conditions holds:
1. $P^{*}(Y|do(X))$ is trivially transportable.
2. There exists a set of covariates $Z$ (possibly affected by $X$) such that $Z$ is $S$-admissible and for which $P^{*}(z|do(x))$ is transportable.
3. There exists a set of covariates, $W$ that satisfy $(X\perp\!\!\!\!\perp Y|W)_{\overline{X(W)}}$ and for which $P^{*}(w|do(x))$ is transportable.
Proof.
1. According to Definition A.4, the causal relationship can be directly estimated from observational data of $\Pi^{*}$.
2. If condition 2 holds, it implies:
$P^{*}(Y|do(X))$ $=P(Y|do(X),S)$ (9)
$=\sum_{Z}P(Y|do(X),Z,S)P(Z|do(X),S)$
$=\sum_{Z}P(Y|do(X),Z)P^{*}(Z|do(X))$
The transportability of $P(Z|do(X))$ reduces $P^{*}(Z|do(X))$ to a star-free expression and therefore $P^{*}(Y|do(X))$ is transportable. 3. If condition 3 holds, it implies:
$P^{*}(Y|do(X))$ $=P(Y|do(X),S)$ (10)
$=\sum_{W}P(Y|do(X),W,S)P(W|do(X),S)$
$=\sum_{W}P(Y|W,S)P^{*}(W|do(X))$
According to Rule 3 of Definition A.2, the transportability of $P^{*}(W|do(X))$ would render $P(W|do(X),S)$ to a star-free expression and render $P^{*}(Y|do(X))$ transportable. This ends the proof. ∎
We are now prepared to present the proof for Theorem 4.5 in the main text:
Theorem A.6.
The formulation of the causal mechanisms $P^{*}(Y|do(X))$ in Figure 2(a) can be written as $\sum_{Z_{var}}P(Y|do(X),Z_{var})P^{*}(Z_{var})$ (see Equation 15 in Appendix A for details). Where the $P(Y|do(X),Z_{var})$ is the invariance causal mechanism that doesn’t depend on the environment $E$. However, without the observation of $Z_{var}$ and $Z_{inv}$, the causal mechanism $P^{*}(Y|do(X))$ in Figure 2(a) is not transportable from $\Pi$ to $\Pi^{*}$.
Proof.
Suppose that the causal mechanism $P^{*}(Y|do(X))$ in Figure 2(a) is transportable from $\Pi$ to $\Pi^{*}$, we aim to prove our conclusion by reduction to absurdity. The causal mechanism we would like to estimate is:
$P^{*}(Y|do(X))$ $=P(Y|do(X),E=e^{*})$ (11)
$=\sum_{Z_{inv}}\sum_{Z_{var}}P(Y|do(X),Z_{var},Z_{inv},E=e^{*})P(Z_{var}|do(x),E=e^{*})P(Z_{inv}|do(x),E=e^{*})$
From the Rule 1 of Definition A.2, we have
$P(Y|do(X),Z_{var},Z_{inv},E=e^{*})=P(Y|do(X),Z_{var},Z_{inv}),$ (12)
because $Z_{var}$ satisfies $(Y\perp\!\!\!\!\perp E|Z_{var},X)$ in $D_{\overline{X}}$. Therefore, according to Definition 4.4, the $Z_{var}$ $E$-admissible with respect to the causal effect of $X$ on $Y$. According to Definition A.1, $Z_{var}\rightarrow Y\leftarrow X$ is a collision node, then:
$P(Z_{var}|do(x),E=e^{*})=P^{*}(Z_{var}),$ (13)
and similarly for $Z_{inv}\rightarrow Y\leftarrow X$ and $Z_{inv}\rightarrow Y\leftarrow Z_{var}\leftarrow E$:
$P(Z_{inv}|do(x),E=e^{*})=P(Z_{inv}).$ (14)
Therefore, the Equation 11 can be rewrite as:
$P^{*}(Y|do(X))=\sum_{Z_{inv}}\sum_{Z_{inv}}P(Y|do(X),Z_{var},Z_{inv})P(Z_{inv})P^{*}(Z_{var})$ (15)
However, in Figure 2(a), the $Z_{var}$ and $Z_{inv}$ is unobservable, making the causal effect $P^{*}(Y|do(X))$ non-transportable. This ends the proof. ∎
Appendix B Further analysis on the invariant causal mechanism
To give a more comprehensive understanding of the invariant causal mechanism, we further provide a detailed analysis as follows: From Equation 15, the distribution of latent variable $Z_{inv}$ is invariant across domains, while the distribution of $Z_{var}$ is dependent on the environment.
Without loss of generality, let $Z=[Z_{inv},Z_{var}]$, $\|Z_{inv}\|=\|Z_{var}\|=1$, the prediction of CLIP can be written as:
$P(Y|do(X),Z_{var},Z_{inv})$ $=\frac{\exp(Z^{T}Z^{Y})}{\sum_{c\in C}\exp(Z^{T}Z^{c})},$ (16)
$=\frac{\exp(Z^{T}_{inv}Z^{Y}_{inv})\cdot\exp(Z^{T}_{var}Z^{Y}_{var})}{\sum_{c\in C}\exp(Z^{T}_{inv}Z^{c}_{inv})\exp(Z^{T}_{var}Z_{var}^{c})},$ (17)
if we divide $\exp(Z^{T}_{var}Z^{Y}_{var})$ both above and below:
$P(Y|do(X),Z_{var},Z_{inv})=\frac{\exp(Z^{T}_{inv}Z^{Y}_{inv})}{\sum_{c\in C}\exp(Z^{T}_{inv}Z^{c}_{inv})\exp(Z^{T}_{var}(Z_{var}^{c}-Z^{Y}_{var}))}.$ (18)
And $\exp(Z^{T}_{var}(Z_{var}^{c}-Z^{Y}_{var}))=\exp(\|Z_{var}\|\cdot\|Z_{var}^{c}-Z^{Y}_{var}\|\cdot\cos\alpha)\leq\exp(\|Z_{var}^{c}-Z^{Y}_{var}\|)$, where $\alpha$ is the angle between $Z_{var}$ and $Z_{var}^{c}-Z^{Y}_{var}$. Therefore:
$\frac{\exp(Z^{T}_{inv}Z^{Y}_{inv})}{\sum_{c\in C}\exp(Z^{T}_{inv}Z^{c}_{inv})\exp(Z^{T}_{var}(Z_{var}^{c}-Z^{Y}_{var}))}\geq\frac{\exp(Z^{T}_{inv}Z^{Y}_{inv})}{\sum_{c\in C}\exp(Z^{T}_{inv}Z^{c}_{inv})\exp(\|Z_{var}^{c}-Z^{Y}_{var}\|)}$ (19)
According to Hölder’s inequality:
$\sum_{c\in C}\exp(Z^{T}_{inv}Z^{c}_{inv})\exp(\|Z_{var}^{c}-Z^{Y}_{var}\|)\leq\sum_{c\in C}\exp(Z^{T}_{inv}Z^{c}_{inv})\sum_{c}\exp(\|Z_{var}^{c}-Z^{Y}_{var}\|).$ (20)
Without loss of generality, suppose $\exists M>0$ is a constant, $s.t.\sum_{c\in C}\exp(\|Z_{var}^{c}-Z^{Y}_{var}\|)=M$. Then the following inequality holds:
$\frac{\exp(Z^{T}_{inv}Z^{Y}_{inv})}{\sum_{c\in C}\exp(Z^{T}_{inv}Z^{c}_{inv})\exp(Z^{T}_{var}(Z_{var}^{c}-Z^{Y}_{var}))}$ $\geq\frac{\exp(Z^{T}_{inv}Z^{Y}_{inv})}{\sum_{c\in C}\exp(Z^{T}_{inv}Z^{c}_{inv})\exp(\|Z_{var}^{c}-Z^{Y}_{var}\|)}$ (21)
$\geq\frac{\exp(Z^{T}_{inv}Z^{Y}_{inv})}{M\sum_{c\in C}\exp(Z^{T}_{inv}Z^{c}_{inv})}$ (22)
Theoretically, maximizing the RHS of the above inequality is a sufficient condition for maximizing the LHS, which is the prediction of CLIP. While the RHS is the invariant prediction of Equation 6 of the original paper, i.e. the invariant causal mechanism.
Empirically, the text prompts are fixed across different domains, e.g. ”A photo of [cls]”. Therefore, the $Z_{var}^{c}$ is fixed across different domains, while $Z_{var}$ is the variance part of image embedding and changes across domains. To make an invariant prediction, the $Z_{var}$ part of the embedding of prompt of each class should be the same, $Z_{var}^{c}=Z_{var}^{Y},\forall c\in C$, therefore, $\exp(Z^{T}_{var}(Z_{var}^{c}-Z^{Y}_{var}))=1$. The invariant prediction can be written as:
$P(Y|do(X),Z_{inv})=\frac{\exp(Z^{T}_{inv}Z^{Y}_{inv})}{\sum_{C}\exp(Z^{T}_{inv}Z^{c}_{inv})}$ (23)
Appendix C Proof for Proposition 4.6
Building upon prior works (Roeder et al., 2021; Ahuja et al., 2022b; Hyvarinen & Morioka, 2016; Zimmermann et al., 2022), we offer proof demonstrating that CLIP identifies latent factors up to an invertible linear transformation, and we also reuse some of their proof techniques.
Proposition C.1.
For an injective data generating process $\boldsymbol{x}=g(Z)$, suppose there exists an ideal encoder $f_{\theta^{\prime}}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}$, such that $f_{\theta^{\prime}}(\boldsymbol{x})=Z$. If for all environment $E_{all}$, the encoder is trained optimal, and there exists a predictor for each environment, such that $p_{\theta}(y^{e}|\boldsymbol{x}^{e};W_{\phi_{e}})=p_{\theta^{\prime}}(y^{e}|\boldsymbol{x}^{e};W^{e}_{\phi^{\prime}_{e}}),\forall e\in E_{all}$. Then the encoder $f_{\theta}(\boldsymbol{x})$ identifies the latent factor $Z$ up to an invertible linear transformation, i.e. $f_{\theta}(\boldsymbol{x})=AZ$.
Proof.
We proceed by constructing the invertible linear transformations to satisfy $f_{\theta}(\boldsymbol{x})=AZ$, where $A$ is an invertible matrix. First, we assume that for any $(\theta,\theta^{\prime})$ for which it holds that $p_{\theta}=p_{\theta^{\prime}}$, and for any given $\boldsymbol{x}$, we assume that by repeated sampling class $C$ from the observation data and picking $y_{A},y_{B}\in C$, we can construct a set of distinct tuples ${(y_{A}^{i},y_{B}^{i})},i=1,2,...,d$, such that the matrices $L$ and $L^{\prime}$ are invertible, where $L$ consists of columns $W_{\phi_{e}}(y_{A}^{i})-W_{\phi_{e}}(y_{B}^{i})$ and $L^{\prime}$ consists of columns $W^{e}_{\phi^{\prime}_{e}}(y_{A}^{i})-W^{e}_{\phi^{\prime}_{e}}(y_{B}^{i})$. The likelihood ratios for these points are equal:
$\frac{p_{\theta}(y_{A}|\boldsymbol{x}^{e},C)}{p_{\theta}(y_{B}|\boldsymbol{x}^{e},C)}=\frac{p_{\theta}^{\prime}(y_{A}|\boldsymbol{x}^{e},C)}{p_{\theta}^{\prime}(y_{B}|\boldsymbol{x}^{e},C)}.$ (24)
Substituting our model definition from Equation 3, we find:
$\frac{exp(f_{\theta}(\boldsymbol{x})^{T}W_{\phi_{e}}(y_{A}^{i}))}{exp(f_{\theta}(\boldsymbol{x})^{T}W_{\phi_{e}}(y_{B}^{i}))}=\frac{exp(f_{\theta}^{\prime}(\boldsymbol{x})^{T}W^{e}_{\phi^{\prime}_{e}}(y_{A}^{i}))}{exp(f_{\theta}^{\prime}(\boldsymbol{x})^{T}W^{e}_{\phi^{\prime}_{e}}(y_{B}^{i}))}.$ (25)
Taking the logarithm of both sides, this simplifies to:
$(W_{\phi_{e}}(y_{A}^{i})-W_{\phi_{e}}(y_{B}^{i}))^{T}f_{\theta}(\boldsymbol{x})=(W^{e}_{\phi^{\prime}_{e}}(y_{A}^{i})-W^{e}_{\phi^{\prime}_{e}}(y_{B}^{i}))^{T}f_{\theta}^{\prime}(\boldsymbol{x}).$ (26)
By repeating Equation 26 $d$ times and by the diversity condition noted above, the resulting difference vectors are linearly independent. We collect these vectors together as the columns of $d\times d$-dimensional matrices $L$ and $L^{\prime}$, forming the following system of $d$ linear equations:
$f_{\theta}(\boldsymbol{x})=(L^{\prime}L^{-1})^{T}f_{\theta}^{\prime}(\boldsymbol{x}).$ (27)
Therefore, $f_{\theta}(\boldsymbol{x})=Af_{\theta}^{\prime}(\boldsymbol{x})$ where $A=L^{\prime}L^{-1}$ is invertible. ∎
Appendix D Proof for Theorem 4.7
Theorem D.1.
Consider observational data generated from an injective mapping $\boldsymbol{x}=g(Z)$, where the latent factors $Z$ follow a distribution $P_{Z}$ supported on $\mathcal{Z}$. If a $do$-intervention is applied to the latent factor $z_{i}$, with the distribution of other latent factors denoted as $P^{do(z_{i})}_{Z}$, the resulting distribution of the interventional data is denoted as $P^{do(z_{i})}_{X}$. Suppose that an encoder $f_{\theta}$ satisfies Proposition 4.6, and the output of the $k^{th}$ component of the encoder is required to take a fixed value $f_{\theta}(\boldsymbol{x})[k]=z^{\dagger}$ for all $x\in\mathcal{X}^{do(z_{i})}$. Then, the encoder identifies the intervened latent $z_{i}$ up to a shift and scaling, i.e., $f_{\theta}(\boldsymbol{x})[k]=a_{i}z_{i}+b_{i}$, where $a_{i}\in\mathbb{R}$ and $b_{i}\in\mathbb{R}$.
Proof.
The encoder $f_{\theta}$ satisfies Proposition 4.6 implies that $f_{\theta}(\boldsymbol{x})=AZ$. The $k$-th element of $f_{\theta}(\boldsymbol{x})[k]=A_{k,:}Z,\forall z\in\mathcal{Z}$. Considering the support of the distribution of the unintervened latent $P^{do(z_{i})}_{Z}$ is a non-empty subset. We separately write the intervened latent as $Z=[z_{i},Z_{i:}]$, where $z_{i}$ is the intervened latent and $Z_{i:}$ is the set of other latent factors other than $z_{i}$. Then, $f_{\theta}(\boldsymbol{x})[k]$ can be written as:
$f_{\theta}(\boldsymbol{x})[k]=A_{k,i}\cdot z_{i}+A_{k,i:}Z_{i:}+b_{i},$ (28)
where $A_{k,i}$ represents the $i$-th element of the $k$-th row, and $A_{k,i:}$ denotes the $k$-th row excluding the $i$-th element. If the output of the $k$-th component of the encoder to have a fixed value of $z^{\dagger}$, then setting $f_{\theta}(\boldsymbol{x})[k]=z^{\dagger}$ yields the following:
$z^{\dagger}=A_{k,i}\cdot z_{i}+A_{k,i:}Z_{i:}+b_{i}.$ (29)
Consider another data point $Z^{\prime}$ from the interventional distribution $P^{do(z_{i})}Z$. We can express $Z^{\prime}$ as $Z^{\prime}=[z_{i},Z^{\prime}_{i:}]$. The $Z^{\prime}$ is nearly identical to $Z$, except for $Z^{\prime}_{i:}$, which is obtained by adding $\delta\boldsymbol{e}_{j}$ to $Z_{i:}$. Here, $\delta$ is a small positive value, and $\boldsymbol{e}_{j}$ represents a one-hot vector with the $j$-th element as one and the others as zero.
$z^{\dagger}=A_{k,i}\cdot z_{i}+A_{k,i:}Z^{\prime}_{i:}.$ (30)
If we use Equation 29 minus Equation 30, the results are as followed:
$A_{k,i:}Z^{\prime}_{i:}$ $=A_{k,i:}Z_{i:}$ (31)
$A_{k,i:}(Z^{\prime}_{i:}-Z_{i:})$ $=0$
$\delta A_{k,i:}\boldsymbol{e}_{j}$ $=0.$
This implies the $j$-th element in $A_{k,i:}$ is zero. Iterative apply $j$ to the $Z_{i:}$ can proof that $A_{k,i:}$ is zero, which makes $f_{\theta}(\boldsymbol{x})[k]=A_{k,i}z_{i}$. This ends the proof. ∎
Appendix E Proof for Theorem 4.8
Before we proceed to the proof of Theorem 4.8, we first give the generalization bound theorem in (Mohri et al., 2018) and illustrate Massart’s lemma, i.e. Lemma E.2.
Theorem E.1.
Given the data distribution $\mathcal{D}$, then for any hypothesis $h$ in the finite set $H\in\mathbb{R}^{m}$ and $1\textgreater\delta\textgreater 0$, with probability at least $1-\delta$, it holds that:
$\begin{split}R^{OOD}(H,\mathcal{D})\leq\frac{1}{m}\sum_{i=1}^{m}\hat{E}(h_{i})+\mathfrak{R}_{m}(H)+\sqrt{\frac{\log(1/\delta)}{2m}},\end{split}$ (32)
where $\mathfrak{R}_{m}(H)$ represents Rademacher complexity.
Lemma E.2.
Let $A\subset\mathbb{R}^{m}$ be a finite set, with $r=max_{\boldsymbol{x\in A}}\|\boldsymbol{x}\|_{2}$, then the following holds:
$\mathbb{E}_{\sigma}\Bigg{[}\frac{1}{m}\sup_{\boldsymbol{x}\in\mathcal{A}}\sum_{i=1}^{m}\sigma_{i}\boldsymbol{x}_{i}\Bigg{]}\leq\frac{r\sqrt{2\log|A|}}{m},$ (33)
where $\sigma_{i}$s are independent uniform random variables taking values in $\{-1,+1\}$ and $x_{1},...,x_{m}$ are the components of vector $\boldsymbol{x}$.
Proof.
The result follows immediately from the bound on the expectation of a maximum given by Corollary D.11 from (Mohri et al., 2018) since the random variables $\sigma_{i}\boldsymbol{x}_{i}$ are independent and each $\sigma_{i}\boldsymbol{x}_{i}$ takes values in $[-|x_{i}|,|x_{i}|]$ with $\sqrt{\sum_{i=1}^{m}x_{i}^{2}}\leq r^{2}$. ∎
Note that the left side of Equation 33 is the formulation of Rademacher complexity, Lemma E.2 provides an upper bound of Rademacher complexity for arbitrary space. Using this result, we can now demonstrate Theorem 4.8.
Theorem E.3.
For any data distribution $\mathcal{D}$, hypothesis spaces $H_{Z},H_{\hat{Z}}\subset\mathbb{R}^{m}$, the generalization error of hypothesis $W_{Z}\in H_{Z}$ and hypothesis $W_{\hat{Z}}\in H_{\hat{Z}}$ satisfies the following inequality:
$R^{OOD}(H_{Z},\mathcal{D})\leq R^{OOD}(H_{\hat{Z}},\mathcal{D}).$ (34)
Proof.
Due to the causality-based identification process of our method, the dimension of hypothesis space $H_{Z}$ and $H_{\hat{Z}}$ satisfies inequality:
$|H_{Z}|\leq|H_{\hat{Z}}|.$ (35)
Further, the upper bound of Rademacher complexity of hypothesis space $H_{Z}$ and $H_{\hat{Z}}$ satisfies:
$\frac{r\sqrt{2\log|H_{Z}|}}{m}\leq\frac{r\sqrt{2\log|H_{\hat{Z}}|}}{m}.$ (36)
The result indicates that the Rademacher complexity satisfies:
$\mathfrak{R}(H_{Z},\mathcal{D})\leq\mathfrak{R}(H_{\hat{Z}},\mathcal{D}).$ (37)
According to (Mohri et al., 2018), the upper bound of generalization error satisfies Theorem E.1. For the generalization bound of CLIP-ICM and CLIP only differs in the Rademacher complexity item, we can dedicate from the above illustration that CLIP-ICM has a lower generalization bound than CLIP:
$R^{OOD}(H_{Z},\mathcal{D})\leq R^{OOD}(H_{\hat{Z}},\mathcal{D}).$ (38)
The above result indicates that the OOD generalization ability of our proposed method is better compared with CLIP. This ends our proof. ∎
E.1 Pseudo Code
The pseodo code of CLIP-ICM is illustrated in Algorithm 1.
Algorithm 1 CLIP-ICM
0:  CLIP image encoder $f_{I}$ and images from training domain $\boldsymbol{x}\in\mathcal{D}$, $\mathcal{D}=\{\mathcal{D}^{e}\}_{e\in E_{tr}}$. $K$ types of data augmentation $\mathcal{A}=\{\alpha_{k}\}_{k=1}^{K}$. Hyperparameter $\lambda$ and $D_{inv}$.
1:  Initialize the container $\hat{Z}=[\quad]$ for original representation
2:  Initialize the container $\hat{Z}^{\prime}=[\quad]$ for intervened representation
3:  for $\boldsymbol{x}$ in $\mathcal{D}$ do
4:    Sample random augmentation $\alpha_{k}$ from $K$
5:    Get the original image representation $\hat{z}=f_{I}(\boldsymbol{x})$
6:    Get $\boldsymbol{x}^{do(k)}=\alpha_{k}(\boldsymbol{x})$
7:    Get the intervened representation $\hat{z}^{\prime}=f_{I}(\boldsymbol{x}^{do(k)})$
8:    Append $\hat{z}$ into $\hat{Z}$
9:    Append $\hat{z}^{\prime}$ into $\hat{Z}^{\prime}$
10:  end for
11:  Get the correlation matrix $Corr=\hat{Z}^{T}\hat{Z}-\lambda\cdot(\hat{Z}-\hat{Z}^{\prime})^{T}(\hat{Z}-\hat{Z}^{\prime})$. $\hat{Z}\in\mathbb{R}^{N\times d},\hat{Z}^{\prime}\in\mathbb{R}^{N\times d},Corr\in\mathbb{R}^{d\times d}$
12:  Perform Singular Value Decomposition (SVD) for $Corr$ to get $U,V,\Sigma$
13:  Get the highest $D_{inv}$ singular vector $U_{:D_{inv},:}\in\mathbb{R}^{D_{inv}\times D}$
13:  The target transform $C_{\xi}$ is $U_{:D_{inv},:}$
Appendix F Dataset Details
DOMAINBED includes downloaders and loaders for seven multi-domain image classification tasks:
[…]
For all datasets, we first pool the raw training, validation, and testing images together. For each random seed, we then instantiate random training, validation, and testing splits.
Appendix G Experiments on ImageNet-C
To evaluate the robustness of our proposed CLIP-ICM, we also conduct experiments on the corruption-based dataset ImageNet-C (Hendrycks & Dietterich, 2019). ImageNet-C is an open-source dataset that consists of algorithmically generated corruptions (blur, noise) applied to the ImageNet test set. We report the mean Corruption Error (mCE) to evaluate the robustness of our method. In addition to the mCE metric, we have also compared the clean error rates. Our comparative analysis includes the evaluation of a ResNet50 baseline, CLIP Linear-probe, and CLIP-ICM Linear-probe. It is noteworthy that all these methods utilize the ResNet50 architecture as their backbone network.
Algorithm mCE Clean Error
ResNet-50 Baseline 76.7% 23.85%
CLIP Linear-probe 62.92% 26.70%
CLIP-ICM Linear-probe 51.74% 22.67%
Table 7: The comparison results of mean Corruption Error and Clean Error on ImageNet-C dataset. A lower value of error rate indicates a better robustness performance.
From Table 7, it is clear that our method achieves superior performance on the ImageNet-C dataset compared to the ResNet-50 Baseline. Additionally, it shows an improvement in performance on the original ImageNet dataset as well.
Appendix H More comparison with related works
H.1 Model-based Domain Generalization (Robey et al., 2021)
Model-based Domain Generalization (Robey et al., 2021) also also proposes an SCM, their analysis method is different from ours. They believe that in domain shift, an instance $X^{e}$ is generated by an underlying random variable $X$ and $e$ jointly passing through a domain-transfer model $G(X,e)$. In our causal diagram analysis, we believe that the sample is generated by $Z_{var}$ and $Z_{inv}$ jointly.
MBDG obtains the invariant features by utilizing a pre-trained domain-transfer model to constrain the distance between the features extracted by the feature extractor and the representations generated by the domain-transfer model.
In contrast, our approach learns a linear transformation of the pre-trained representation to obtain features that are invariant to data augmentation. We provide the comparison results between our method and MBDG in Table 2 and Table 3.
H.2 On Calibration and Out-of-domain Generalization (Wald et al., 2021)
From the perspective of SCM analysis, (Wald et al., 2021) proposes a causal graph that categorizes features into $X_{causal}$, $X_{ac-spurious}$, and $X_{ac-non-spurious}$. This categorization is based on the understanding that certain variables have a direct causal influence on the outcome $Y$. In contrast, our proposed causal graph incorporates the environmental variable $E$ as an indirect causal pathway to $Y$.
From the methodological standpoint, the article addresses the OOD problem by proposing the need to find a classifier $f$ such that $Y\perp E|f(X)$, which implies that the classifier $f$ is calibrated. Similar to IRM, the article introduces a loss function called CLOvE designed to constrain classifiers across different domains to produce consistent results. Our approach, however, does not require modifications to the backbone model, nor do we necessitate the training of additional classifiers. Instead, we achieve a zero-shot classifier that remains invariant across each domain.
H.3 Subspace Identification Guarantee (Li et al., 2024)
From the perspective of SCM analysis, SIG views the data generation process through 4 kinds of latent factors. In our SCM, $Z_{inv}$ can be regarded as a domain-invariant variable, and $Z_{var}$ can be regarded as a domain-variant variable. Since in our causal diagram, $Y$ represents the description of the picture, therefore, we do not distinguish between label-irrelevant and label-relevant variables in our modeling.
From a methodological perspective, SIG uses an end-to-end, reconstruction-based network, while we use a pre-trained CLIP backbone network. SIG uses a variation-inference method to identify latent factors, while we use an intervention method to identify latent factors.
A comparative analysis is presented to illustrate the performance of our method alongside the SIG’s approach to the Office-Home dataset. Both our method and the SIG’s method utilize the ResNet50 architecture as the backbone network. Similarly, the performance on the DomainNet dataset is compared, with both our method and the SIG’s method employing the ResNet101 backbone network.
H.4 Domain Prompt Learning (Zhang et al., 2022)
(Zhang et al., 2022) introduces Domain Prompt Learning (DPL), a novel approach that improves the domain generalization performance of the CLIP model on various datasets by generating conditional prompts, achieving higher accuracy without fine-tuning the entire model. Compared with our method, also includes training a fully connected network with label information in the objective function. We provide the comparison results below:
Table 8: Comparison between our method and DPL across various datasets.
Method VLCS PACS OfficeHome Terra Avg
CLIP Linear-Probe 78.7 96.4 81.9 60.2 79.3
CLIP + DPL (Zhang et al., 2022) 84.3 97.3 84.2 52.6 79.6
CLIP-ICM Linear-Probe 86.5 97.5 84.6 64.3 83.2
H.5 Self-Supervised Learning with Data Augmentations Provably Isolates Content from Style (Kügelgen et al., 2021)
(Kügelgen et al., 2021) also interpret data augmentations as counterfactuals, and use them to obtain the invariant factors. Our work differs from theirs in three key aspects: the background setting, the research problem, and the implementation method.
Firstly, the problem studied by (Kügelgen et al., 2021) is set within the framework of self-supervised contrastive learning. In their paper, the term ”invariant factor” refers to the invariant parts derived from two different augmented perspectives within a positive sample pair. In contrast, the ”invariant factor” in our paper refers to the latent factors that remain unchanged across different domains. Due to the diversity of augmentation methods employed in self-supervision (Chen et al., 2020), the invariant factors they investigate do not align with the invariant factors we examine.
Secondly, in their paper, they consider an in-distribution problem, where all representations $Z$ belong to the same distribution $Z\sim P(Z)$, and they do not take into account the different distributions across various domains. In our paper, however, we investigate an out-of-distribution problem. We posit that the distribution of $Z_{var}$ is influenced by the environment, meaning that $P(Z_{var})\neq P^{*}(Z_{var})$.
Lastly, their experiments are solely based on analyzing the capabilities of SimCLR to identify latent factors, without proposing any new methods. In contrast, based on our theoretical analysis, we proposed a novel method for identifying invariant representations.
Table 9: Comparison results on DomainNet dataset with ResNet101 backbone.
Algorithm Backbone Clipart Infograph Painting Quickdraw Real Sketch Avg
SIG (Li et al., 2024) R101 72.7 32.0 61.5 20.5 72.4 59.5 53.0
CLIP-ICM Linear-probe R101 73.6 38.2 69.4 19.2 74.1 65.5 56.6
Appendix I Ablation study on dimension of $Z_{inv}$
As we mentioned at Section 4.3, the dimension $D_{inv}$ of $Z_{inv}$ is a hyperparameter. To investigate the influence of $D_{inv}$, we conduct an ablation study regarding the choice of $D_{inv}$. All results are the zero-shot performance of CLIP-ICM, conducted on the VLCS dataset with ViT-B/16 as the backbone, where $D_{inv}<512$. The results are shown in Figure 8.
Figure 8: The accuracy of CLIP-ICM on VLCS with different choice of $D_{inv}$
From Figure 8 we can see that as the $D_{inv}$ increases, the accuracy first increases, and stabilizes, and slowly decreases, and finally drops. Therefore, the optimum dimension of $D_{inv}$ should be around $300$ to $350$ for all domains in VLCS.
Appendix J Ablation study on zero-shot prediction with domain prompt
In experiments from Section 5, we use the standard text prompt template when performing zero-shot prediction, i.e. ”A photo of [cls]” where [cls] is the name of the corresponding class.
To explore the influence of text prompts on zero-shot prediction, we conducted experiments using the PACS, VLCS, and Office-Home datasets. In these experiments, we employed text prompts that incorporate domain information. For datasets where the domain names hold particular significance, such as PACS and Office-Home, we utilized the prompt ’A [domain] of [cls]’. In contrast, for the VLCS dataset, where the domain names do not have specific meanings, we adopted the prompt ’A photo of [cls] in [domain]’. Here, [domain] represents the name of the test domain. The results are illustrated in Table 10, Table 11 and Table 12.
It is important that the objective of domain generalization is to evaluate the performance of a model on unseen domains. Therefore, incorporating domain information, such as (a sketch of [cls]) would not align with the requirements of the domain generalization task. And we only conduct this experiment for exploration.
Table 10: Comparison results of zero-shot performance of CLIP with domain prompt in PACS dataset. $\uparrow$ denotes that the results of CLIP with domain prompt are higher than standard prompt, while $\downarrow$ denotes that the results of CLIP with domain prompt are lower than standard prompt.
Algorithm Photo Art Cartoon Sketch Avg
CLIP 100.0 97.4 99.2 88.1 96.1
CLIP + Domain Prompt 100.0 -- 97.4 -- 99.0 $\downarrow$ 90.2 $\uparrow$ 96.6 $\uparrow$
CLIP-ICM 98.7 99.1 99.9 91.7 97.3
Table 11: Comparison results of zero-shot performance of CLIP with domain prompt in VLCS dataset. $\uparrow$ denotes that the results of CLIP with domain prompt are higher than standard prompt, while $\downarrow$ denotes that the results of CLIP with domain prompt are lower than standard prompt.
Algorithm Caltech101 LabelMe SUN09 VOC2007 Avg
CLIP 99.9 70.1 73.5 86.1 82.4
CLIP + Domain Prompt 99.2 $\downarrow$ 69.7 $\downarrow$ 65.1 $\downarrow$ 80.3 $\downarrow$ 78.5 $\downarrow$
CLIP-ICM 100.0 74.7 74.5 87.1 84.1
Table 12: Comparison results of zero-shot performance of CLIP with domain prompt in Office-Home dataset. $\uparrow$ denotes that the results of CLIP with domain prompt are higher than standard prompt, while $\downarrow$ denotes that the results of CLIP with domain prompt are lower than standard prompt.
Algorithm Art Clipart Product Real Avg
CLIP 83.3 65.3 89.0 89.3 81.7
CLIP + Domain Prompt 82.4 $\downarrow$ 67.0 $\uparrow$ 87.8 $\downarrow$ 88.7 $\downarrow$ 81.4 $\downarrow$
CLIP-ICM 83.7 67.4 90.8 90.1 82.6
From the results in Table 10, we find that only in some domains, incorporating domain information in the prompt template can improve the zero-shot performance, such as Sketch in PACS, and Clipart in Office-Home. However, in most domains, where the domain name doesn’t provide any information, or can’t describe the domain properly, the performance is lower than the original template.
Another important observation is that CLIP-ICM consistently improves the performance of CLIP in most domains, without any domain-related information or other kinds of template design.